** Please use ICE Table for these problems**

1. A substance MX has a Ksp of 1.27x10^-36

a. What is it's molar solubility in pure water? (hint: ICE Table-reminder MX will fall out of your equilibrium expression)

b.What is it's molar solubility in 0.25 MBr2? (hint: ICE Table-but what is M2+ initial)

2.A solution that is 0.200M HF and 0.240 M NaF reacts to produce a mixture with what pH? The pKa of HF is 3.20. (hint: ICE Table and x)

if you can show all the steps with ice table I would really appreciate it!
Answer #1

MX (s) m*lag) + x Lag) ; Ksp = 1.27x10-36 xM xM JR X M of Mx be soluble , then [M+9 =[*-) = xM Ksp = [M+9[*-4 = 1,27 *10-36 oN.B 1: There is no meaning of 0.25 MBr2. It could be 0.25M Br2 or 0.25M MBr2. We expect it was 0.25M MBr2.

N.B 2: As per Chegg guidelines we have to answer the first big answer type question among multiple questions. We expect you can solve the rest part, otherwise reupload the rest portion again.